# Ceresa cycles of bielliptic Picard curves

### v1

@@ 1.4. Previous results
Several examples of nonhyperelliptic curves with torsion Ceresa cycle were recently found, conditional on Beilinson–Bloch type conjectures [7, 20, 4, 5, 30]. Qiu and W. Zhang then gave some unconditional examples [37], including one in genus 3 and one-dimensional families in genus $4$ and $5$. Laterveer also found a two-dimensional family in genus $5$ [29]. Subsequently, Qiu–Zhang found examples of plane quartic curves (dominated by Shimura curves) with trivial automorphism group and torsion Ceresa cycle [38].
The proofs of [37] and [29] rely on the vanishing of the Chow motive $(\mathfrak{h}^{1}(C)^{\otimes 3})^{{\rm Aut}(C)}$, and hence of $\mathfrak{h}^{3}(J)^{{\rm Aut}(C)}$, for certain curves $C$, generalizing the observation that hyperelliptic curves have trivial Ceresa cycle. For our curves $C=C_{t}$ with $t\neq 0$, the motive $\mathfrak{h}^{3}(J)^{{\rm Aut}(C)}=\mathfrak{h}^{3}(J)^{\mu_{6}}$ does not vanish, but (1.2) shows that it is a direct summand of the motive of a curve, hence is still reasonably accessible. It would be interesting to study other families of curves where $\mathfrak{h}^{3}(J)^{{\rm Aut}(C)}$ is “small” in a suitable sense. More generally, one can try to study the nonzero isotypic components of $\kappa(C)$ with respect to the action of ${\rm End}(J)$ on $\mathfrak{h}^{3}(J)$. For example, even though some of the examples in [38] have trivial automorphism group, they all satisfy ${\rm End}(J)\neq{\rm\mathbb{Z}}$.
Finally, Eskandari and Murty showed that infinitely many Fermat curves have infinite order Ceresa cycle [16], proving a weak version of Corollary 1.2 (where the genus is allowed to grow). If we fix $g$ but allow for curves defined over ${\rm\overline{\mathbb{Q}}}$ (and not necessarily over ${\rm\mathbb{Q}}$), then the analogue of Corollary 1.2 holds for every genus $g\geq 3$ by [44, Theorem 1.3.5].
@@ 1.5. Further questions
For an integer $N\geq 1$, let $Z_{g}(N)$ be the subset of the coarse moduli space of curves $\mathcal{M}_{g}$ over ${\rm\mathbb{Q}}$ consisting of those curves whose canonical Ceresa cycle (with respect to some choice of $(2g-2)$th root of the canonical bundle) is killed by $N$ in the Chow group. What is the geometry of $Z_{g}(N)$? More specifically, what are its positive-dimensional components? Is the union of $Z_{g}(N)$ over all $N$ dense in $\mathcal{M}_{g}$?
@@ 1.6. Acknowledgements
We thank Henri Darmon, David Lilienfeldt, and Padma Srinivasan for helpful conversations and remarks. We thank Adam Logan for numerically testing Theorem 1.1 (while it was still a conjecture), using the forthcoming [15]. We thank Dick Gross, whose question helped motivate us to prove Theorem 1.1. Finally, we thank Tim Dokchitser and Drew Sutherland for helping us compute the central $L$-values mentioned in §6. This research was carried out while the first author was a Research Fellow at St John’s College, University of Cambridge. The second author was funded by the European Research Council (ERC, CurveArithmetic, 101078157).
@@ 1.7. Notation and conventions
  • A variety over a field $k$ is a seperated scheme of finite type over $k$. A variety is nice if it is smooth, projective and geometrically integral.
  • Let $X$ be a smooth and quasi-projective scheme over a Dedekind domain (which might be a field). Then the Chow groups ${\rm CH}^{p}(X)$ of codimension $p$ cycles on $X$ (modulo rational equivalence and with ${\rm\mathbb{Z}}$-coefficients) are defined and come equipped with flat pullbacks, proper pushforwards and an intersection product, see [19, §20.2]. If $Z\subset X$ is a codimension $p$ reduced closed subscheme write $[Z]$ for its class in ${\rm CH}^{p}(X)$.
  • If $X$ is a nice variety over a field $k$ we write ${\rm CH}^{p}(X)_{\mathrm{hom}}\subset{\rm CH}^{p}(X)$ for the classes in the kernel of the $\ell$-adic cycle class map for every prime $\ell$ invertible in $k$.
@@
  • If $A,B\rightarrow S$ are abelian schemes we denote by ${\rm Hom}(A,B)$ the set of $S$-homomorphisms between $A$ and $B$ and ${\rm Hom}^{0}(A,B)={\rm Hom}(A,B)\otimes{\rm\mathbb{Q}}$. We similarly write ${\rm End}(A)={\rm Hom}(A,A)$ and ${\rm End}^{0}(A)={\rm End}(A)\otimes{\rm\mathbb{Q}}$. We write $A^{\vee}\rightarrow S$ for the dual abelian scheme (see [18, Chapter 1, §1]) and write ${\rm Hom}^{\operatorname{sym}}(A,A^{\vee})$ for the subset of all self-dual homomorphisms $A\rightarrow A^{\vee}$.
@@ 2. Ceresa and modified diagonal cycles
We recall the definitions of Ceresa and modified diagonal cycles, and show, by means of an example, a way to verify that $\kappa(C)$ is infinite order, for genus 3 curves over ${\rm\overline{\mathbb{Q}}}$. The latter is based on the methods of [11] and [15], which are in turn based on ideas of Bloch [8]. In this section, and for the rest of the paper, we work over a general (not necessarily algebraically closed) field $k$ of characteristic $\neq 2,3$.
@@ 2.1. Ceresa and modified diagonal cycles
Let $C$ be a (nice) curve of genus $g\geq 2$ over $k$. If $e\in C(k)$, let $\iota_{e}\colon C\hookrightarrow J$ be the Abel-Jacobi map sending $x\mapsto x-e$, and define the Ceresa cycle based at $e$ to be
$\kappa_{e}(C):=[\iota_{e}(C)]-(-1)^{*}[\iota_{e}(C)]\in{\rm CH}^{g-1}(J).$
Lemma 2.1.
The map $e\mapsto\kappa_{e}(C)$ extends to a homomorphism $\alpha_{*}\colon{\rm CH}^{1}(C)\to{\rm CH}^{g-1}(J)$.
Proof.
Let $f\colon C^{2}\to C\times J$ be the morphism $(e,x)\mapsto(e,x-e)$ and let $\alpha^{\prime}=f_{*}[C^{2}]\in{\rm CH}^{2}(C\times J)$. The correspondence $\alpha:=\alpha^{\prime}-(\mathrm{id}\times(-1))^{*}\alpha^{\prime}\in{\rm CH}^{2}(C\times J)$ induces the homomorphism $\alpha_{*}\colon{\rm CH}^{1}(C)\to{\rm CH}^{g-1}(J)$ sending $Z$ to $p_{2*}((p_{1}^{*}Z)\cdot\alpha)$, where $p_{1},p_{2}$ denote the projections from $C\times J$. A computation shows that $\alpha_{*}([e])=\kappa_{e}(C)$ for all $e\in C(k)$. ∎
For a degree 1 divisor class $e\in{\rm CH}^{1}(C)$, we define $\kappa_{e}(C):=\alpha_{*}(e)$. When $(2g-2)e$ is canonical, we call $\kappa_{e}(C)$ a canonical Ceresa cycle and simply write $\kappa(C)$. There are finitely many such $e\in{\rm CH}^{1}(C)$, and they all give the same canonical Ceresa class up to $(2g-2)$-torsion, by Lemma 2.1.
@@
where $\Delta_{123}=\{(c,c,c)\colon c\in C\}$, $\Delta_{12}=\{(c,c,e)\colon c\in C\}$, etc. As before, one can define a cycle $\Delta_{GS,e}$ for any divisor $e$ of degree $1$ on $C$, and if $(2g-2)e$ is a canonical divisor, then we call $\Delta_{GS,e}$ a canonical modified diagonal cycle and write $\Delta_{GS}=\Delta_{GS,e}$. Again this is well-defined up to torsion in the Chow group; see [21, eq. 3.5].
Some useful general facts about these cycles are:
@@
The cycles $\kappa_{e}(C)$ and $\Delta_{GS,e}(C)$ are homologically trivial.
@@
For $\kappa_{e}(C)$, this is immediate from the fact that $(-1)^{*}$ acts trivially on even dimensional cohomology. For $\Delta_{GS,e}(C)$, this is [21, Proposition 3.1]. ∎
Theorem 2.3.
[44, Theorem 1.5.5] $\kappa(C)$ has infinite order if and only if $\Delta_{GS}(C)$ has infinite order.
The following proposition shows that the canonical modified diagonal cycle $\Delta_{GS}(C)$ is “minimal up to torsion” in a certain sense.
@@
If $\Delta_{GS,e}(C)$ is torsion, then $(2g-2)e-K_{C}$ is torsion in $J(k)$, where $K_{C}$ is the canonical divisor class.
@@
This follows from [37, Proposition 2.3(1)], whose proof is valid over any field $k$. ∎
@@ 2.2. A non-vanishing criterion
see [22, §9] for more details concerning this map. The following proposition is based on the methods of [11] and [15].
@@ 2.3. An infinite order example
We now show by example how Proposition 2.5 can be used to verify that $\kappa(C)$ has infinite order. For the rest of the section, let $k={\rm\mathbb{Q}}$, let $C$ be the curve with plane model $y^{3}=x^{4}+x^{2}+1$, and let $\ell=7$. Consider the modified diagonal cycle $\Delta_{GS}$ based at the unique point $\infty$ at infinity; note that $(2g-2)\infty=4\infty$ is canonical.
@@
Since $C$ is a nonhyperelliptic genus $3$ curve, the Abel–Jacobi map $\Sigma\colon{\rm Sym}^{3}(C)\to J$ sending a degree $3$ divisor $D$ to $D-3\infty$ is the blow-up of $J$ along the image of those $D$ with the property that $\dim\operatorname{H}^{0}(C,\mathcal{O}(D))=2$ [33, Theorem 2.3]. By Riemann-Roch, these are exactly the degree $3$ divisors linearly equivalent to $4\infty-x$ for some unique $x\in C$. We conclude that $\Sigma$ is the blow-up of $J$ along the closed subvariety $\{\infty-x\mid x\in C\}\simeq C$. The lemma now follows from the description of the cohomology of a blow-up [1, VII, Theoreme 8.1.1]. ∎
@@
The class $\Delta_{GS}$ has infinite order.
@@
Suppose, for the sake of contradiction, that $\Delta_{GS}$ has finite order. Then its Abel-Jacobi image ${\rm AJ}_{7}(\Delta_{GS})\in\operatorname{H}^{1}({\rm Gal}_{{\rm\mathbb{Q}}},\operatorname{H}^{3}(C^{3}_{\bar{\rm\mathbb{Q}}},{\rm\mathbb{Z}}_{7}))$ has finite order. By Lemma 2.6 it is nonzero. Let $h\colon C^{3}\to{\rm Sym}^{3}(C)$ be the quotient map and let $\tilde{\Delta}_{GS}=h_{*}\Delta_{GS}$. Use the same notation as before for the Abel-Jacobi map ${\rm AJ}_{\ell}\colon{\rm CH}^{2}({\rm Sym}^{3}(C))_{\mathrm{hom}}\to\operatorname{H}^{1}({\rm Gal}_{\rm\mathbb{Q}},V).$ Since $h^{*}\tilde{\Delta}_{GS}=6\Delta_{GS}$, we conclude that ${\rm AJ}_{\ell}(\tilde{\Delta}_{GS})$ is also nonzero and torsion. On the other hand, we may use Magma to compute the characteristic polynomial of a Frobenius element acting on $\operatorname{H}^{1}(C_{{\rm\overline{\mathbb{Q}}}},{\rm\mathbb{Z}}_{7}(1))$, and hence on $V$ as well (via Lemma 2.7). We check that the action of the geometric Frobenius $\mathrm{Fr}_{11}$ at $11$ on $V$ satisfies
@@
Thus $\operatorname{H}^{1}({\rm Gal}_{\rm\mathbb{Q}},V)$ is torsion-free by [43, 6.1], which is a contradiction. ∎
@@
The Ceresa cycle $\kappa(C)$ of the curve $C\colon y^{3}=x^{4}+x^{2}+1$ has infinite order.

### v2

@@ 1.4. Previous results
Several examples of nonhyperelliptic curves with torsion Ceresa cycle were recently found, conditional on Beilinson–Bloch type conjectures [7, 23, 4, 35]. Qiu and W. Zhang then gave some unconditional examples [42], including one in genus 3 and one-dimensional families in genus $4$ and $5$. Laterveer also found a two-dimensional family in genus $5$ [34]. Subsequently, Qiu–Zhang found examples of plane quartic curves (dominated by Shimura curves) with trivial automorphism group and torsion Ceresa cycle [43].
The proofs of [42] and [34] rely on the vanishing of the Chow motive $(\mathfrak{h}^{1}(C)^{\otimes 3})^{{\rm Aut}(C)}$, and hence of $\mathfrak{h}^{3}(J)^{{\rm Aut}(C)}$, for certain curves $C$, generalizing the observation that hyperelliptic curves have trivial Ceresa cycle. For our curves $C=C_{t}$ with $t\neq 0,\pm 1$, the motive $\mathfrak{h}^{3}(J)^{{\rm Aut}(C)}=\mathfrak{h}^{3}(J)^{\mu_{6}}$ does not vanish, but (1.3) shows that it is a direct summand of the motive of a curve, hence is still reasonably accessible. It would be interesting to study other families of curves where $\mathfrak{h}^{3}(J)^{{\rm Aut}(C)}$ is “small” in a suitable sense. More generally, one can try to study the nonzero isotypic components of $\kappa(C)$ with respect to the action of ${\rm End}(J)$ on $\mathfrak{h}^{3}(J)$. For example, even though some of the examples in [43] have trivial automorphism group, they all satisfy ${\rm End}(J)\neq{\rm\mathbb{Z}}$.
Aside from the aforementioned [13], Eskandari and Murty have shown that infinitely many Fermat curves have infinite order Ceresa cycle [18], proving a weak version of Corollary 1.2 (where the genus is allowed to grow). If we fix $g$ but allow for curves defined over ${\rm\overline{\mathbb{Q}}}$ (and not necessarily over ${\rm\mathbb{Q}}$), then the analogue of Corollary 1.2 holds for every genus $g\geq 3$ by [49, Theorem 1.3.5].
Prior to [42], Beauville and Schoen showed that the Ceresa cycle of the curve $y^{3}z=x^{4}+xz^{3}$ is torsion in the Griffiths group [5]. It was observed in [35] that their proof applies to the bielliptic Picard curve $C_{0}$ as well. Theorem 1.5 gives the first example of a family of plane quartic curves with this property.
@@ 1.5. Further questions
Let $V_{g}$ be the subset of the coarse moduli space of curves $\mathcal{M}_{g}$ over ${\rm\mathbb{Q}}$ consisting of those (geometric isomorphism classes of) curves $C$ such that $\kappa(C)\in{\rm CH}_{1}(J)\otimes{\rm\mathbb{Q}}$ vanishes. This is a countable union of closed algebraic subvarieties [33, Lemma 5.1]. Can one describe the positive dimensional components of the torsion locus $V_{g}$? Is $V_{g}$ Zariski dense in $\mathcal{M}_{g}$? After the current paper was posted, several papers have appeared that study this torsion locus and its variants; see [22, 33, 27].
@@ 1.6. Acknowledgements
We thank Henri Darmon, David Lilienfeldt, and Padma Srinivasan for helpful conversations and remarks. We thank Adam Logan for numerically testing Theorem 1.1 (while it was still a conjecture), using the forthcoming [17]. We thank Dick Gross, whose question helped motivate us to prove Theorem 1.1. We thank Tim Dokchitser and Drew Sutherland for helping us compute the central $L$-values mentioned in §6. Finally, we thank the referees for their careful reading of the paper. This research was carried out while the first author was a Research Fellow at St John’s College, University of Cambridge. The second author was funded by the European Research Council (ERC, CurveArithmetic, 101078157).
@@ 1.7. Notation and conventions
  • A variety over a field $k$ is a separated scheme of finite type over $k$. A curve is a smooth, projective and geometrically integral variety of dimension $1$.
  • Let $X$ be a smooth and quasi-projective scheme over a Dedekind domain (which might be a field). Then the Chow groups ${\rm CH}^{p}(X)$ of codimension $p$ cycles on $X$ (modulo rational equivalence and with ${\rm\mathbb{Z}}$-coefficients) are defined and come equipped with flat pullbacks, proper pushforwards and an intersection product, see [21, §20.2]. If $Z\subset X$ is a codimension $p$ reduced closed subscheme write $[Z]$ for its class in ${\rm CH}^{p}(X)$.
  • If $X$ is a smooth projective variety over a field $k$ we write ${\rm CH}^{p}(X)_{\mathrm{hom}}\subset{\rm CH}^{p}(X)$ for the classes in the kernel of the $\ell$-adic cycle class map for every prime $\ell$ invertible in $k$.
@@
  • If $A,B\rightarrow S$ are abelian schemes we denote by ${\rm Hom}(A,B)$ the set of $S$-homomorphisms between $A$ and $B$ and ${\rm Hom}^{0}(A,B)={\rm Hom}(A,B)\otimes{\rm\mathbb{Q}}$. We similarly write ${\rm End}(A)={\rm Hom}(A,A)$ and ${\rm End}^{0}(A)={\rm End}(A)\otimes{\rm\mathbb{Q}}$. We write $A^{\vee}\rightarrow S$ for the dual abelian scheme (see [20, Chapter 1, §1]) and write ${\rm Hom}^{\operatorname{sym}}(A,A^{\vee})$ for the subset of all self-dual homomorphisms $A\rightarrow A^{\vee}$.
@@ 2. Ceresa and modified diagonal cycles
We recall the definitions of Ceresa and modified diagonal cycles, and show, by means of an example, a way to verify that $\kappa_{e}(C)$ is infinite order, for genus 3 curves over ${\rm\overline{\mathbb{Q}}}$. The latter is based on the methods of [11] and [17], which are in turn based on ideas of Bloch [8]. In this section, and for the rest of the paper, we work over a general (not necessarily algebraically closed) field $k$ of characteristic $\neq 2,3$.
@@ 2.1. Ceresa and modified diagonal cycles
Let $C/k$ be a curve of genus $g\geq 2$ with Jacobian $J$. Let $e$ be a degree-$1$ divisor on $C$ and let $\iota_{e}\colon C\rightarrow J$ be the Abel–Jacobi map based at $e$. We define the Ceresa cycle based at $e$ to be
$\kappa_{e}(C):=[\iota_{e}(C)]-(-1)_{*}[\iota_{e}(C)]\in{\rm CH}_{1}(J).$
@@
where $\Delta_{123}=\{(c,c,c)\colon c\in C\}$, $\Delta_{12}=\{(c,c,e)\colon c\in C\}$, and so on. For a general degree $1$ divisor $e=\sum_{i}n_{i}[p_{i}]$ on $C$, we define $\Delta_{GS,e}(C)=\sum_{i}n_{i}\Delta_{GS,p_{i}}\in{\rm CH}^{2}(C^{3})$.
We collect some useful facts about the Ceresa and modified diagonal cycles.
@@
The cycle classes $\kappa_{e}(C)$ and $\Delta_{GS,e}(C)$ are homologically trivial.
@@
For $\kappa_{e}(C)$, this is immediate from the fact that $(-1)^{*}$ acts trivially on even dimensional cohomology. For $\Delta_{GS,e}(C)$, this is [24, Proposition 3.1]. ∎
@@
Suppose that either $\kappa_{e}(C)\in{\rm CH}_{1}(J)$ or $\Delta_{GS,e}(C)\in{\rm CH}^{2}(C^{3})$ is torsion. Then $(2g-2)e-K_{C}$ is torsion in ${\rm CH}_{0}(C)$, where $K_{C}$ is the canonical divisor class.
@@
If $\kappa_{e}(C)$ is torsion, then this follows from [33, Lemma 2.10], but we give an alternative proof here. If $D$ is a divisor class of degree $g-1$ on $C$, let $\Theta_{D}$ denote the image of the map ${\rm Sym}^{g-1}(C)\rightarrow J$ defined by $x\mapsto[x]-D$. If $\kappa_{e}(C)$ is torsion, then $[\iota_{e}(C)]=(-1)_{*}[\iota_{e}(C)]$ in ${\rm CH}_{1}(J)\otimes{\rm\mathbb{Q}}$. Taking the $(g-1)$-fold Pontryagin product of this equality and dividing by $(g-1)!$, we get $[\Theta_{(g-1)e}]=(-1)_{*}[\Theta_{(g-1)e}]$ in ${\rm CH}_{1}(J)\otimes{{\rm\mathbb{Q}}}$. By Riemann–Roch, $(-1)_{*}[\Theta_{(g-1)e}]=[\Theta_{K_{C}-(g-1)e}]$. Since $\Theta_{(g-1)e}$ defines a principal polarization, the map $\varphi\colon J(k)\rightarrow{\rm CH}^{1}(J)_{\hom}$ defined by $x\mapsto[\Theta_{(g-1)e+x}]-[\Theta_{(g-1)e}]$ is an isomorphism. Since
$\varphi(K_{C}-(2g-2)e)=[\Theta_{K_{C}-(g-1)e}]-[\Theta_{(g-1)e}]=(-1)_{*}[\Theta_{(g-1)e}]-[\Theta_{(g-1)e}]$
is torsion in ${\rm CH}^{1}(J)_{\hom}$, the divisor class $K_{C}-(2g-2)e$ is torsion.
If $\Delta_{GS,e}(C)$ is torsion, this follows from [42, Proposition 2.3.2(1)] (where it is assumed that $k={\rm\mathbb{C}}$, but whose proof is valid over any field $k$). ∎
Lemma 2.3.
If $e,e^{\prime}$ are degree-$1$ divisors on $C$ such that $e-e^{\prime}$ is torsion in ${\rm CH}_{0}(C)$, then $\kappa_{e}(C)-\kappa_{e^{\prime}}(C)\in{\rm CH}_{1}(J)$ and $\Delta_{GS,e}(C)-\Delta_{GS,e^{\prime}}(C)\in{\rm CH}^{2}(C^{3})$ are both torsion.
Proof.
The fact that $\kappa_{e}(C)-\kappa_{e^{\prime}}(C)$ is torsion follows from [33, Lemma 2.11], and the fact that $\Delta_{GS,e}(C)-\Delta_{GS,e^{\prime}}(C)$ is torsion follows from [24, Proposition 3.6]. ∎
If $k$ is algebraically closed, choose a degree-$1$ divisor $e$ such that $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)\otimes{{\rm\mathbb{Q}}}$ and define $\kappa(C)$ to be the image of $\kappa_{e}(C)$ in ${\rm CH}_{1}(J)\otimes{{\rm\mathbb{Q}}}$, and $\Delta_{GS}(C)$ to be the image of $\Delta_{GS,e}(C)$ in ${\rm CH}^{2}(C^{3})\otimes{\rm\mathbb{Q}}$. Lemma 2.3 shows that $\kappa(C)$ is independent of the choice of $e$, and Proposition 2.2 shows that $\kappa(C)=0$ if and only if $\kappa_{e}(C)$ is torsion for some degree-$1$ divisor $e$. Similar remarks apply to $\Delta_{GS}(C)$.
Theorem 2.4 (Shou-Wu Zhang).
For every degree-$1$ divisor $e$ on $C$, $\kappa_{e}(C)$ is torsion if and only if $\Delta_{GS,e}(C)$ is torsion.
Proof.
By Proposition 2.2, we may assume $(2g-2)e=K_{C}$ in ${\rm CH}_{0}(C)\otimes{\rm\mathbb{Q}}$. The result then follows from [49, Theorem 1.5.5]. ∎
@@ 2.2. A non-vanishing criterion
see [25, §9] for more details concerning this map. The following proposition is based on the methods of [11] and [17].
@@ 2.3. An infinite order example
We now show by example how Proposition 2.5 can be used to verify that $\kappa_{e}(C)$ has infinite order. For the rest of the section, let $k={\rm\mathbb{Q}}$, let $C$ be the curve with plane model $y^{3}=x^{4}+x^{2}+1$, and let $\ell=7$. Consider the modified diagonal cycle $\Delta_{GS,\infty}$ based at the unique point $\infty$ at infinity; note that $(2g-2)\infty=4\infty$ is canonical.
@@
Since $C$ is a nonhyperelliptic genus $3$ curve, the Abel–Jacobi map $\Sigma\colon{\rm Sym}^{3}(C)\to J$ sending a degree $3$ divisor $D$ to $D-3\infty$ is the blow-up of $J$ along the image of those $D$ with the property that $\dim\operatorname{H}^{0}(C,\mathcal{O}(D))=2$ [38, Theorem 2.3]. By Riemann-Roch, these are exactly the degree-$3$ divisors linearly equivalent to $4\infty-x$ for some unique $x\in C$. We conclude that $\Sigma$ is the blow-up of $J$ along the closed subvariety $\{\infty-x\mid x\in C\}\simeq C$. The lemma now follows from the description of the cohomology of a blow-up [1, VII, Théorème 8.1.1]. ∎
@@
The class $\Delta_{GS,\infty}$ has infinite order.
@@
Suppose, for the sake of contradiction, that $\Delta_{GS,\infty}$ has finite order. Then its Abel-Jacobi image ${\rm AJ}_{7}(\Delta_{GS,\infty})\in\operatorname{H}^{1}({\rm Gal}_{{\rm\mathbb{Q}}},\operatorname{H}^{3}(C^{3}_{\bar{\rm\mathbb{Q}}},{\rm\mathbb{Z}}_{7}))$ has finite order. By Lemma 2.6 it is nonzero. Let $h\colon C^{3}\to{\rm Sym}^{3}(C)$ be the quotient map and let $\tilde{\Delta}_{GS}=h_{*}\Delta_{GS,\infty}$. Use the same notation as before for the Abel-Jacobi map ${\rm AJ}_{\ell}\colon{\rm CH}^{2}({\rm Sym}^{3}(C))_{\mathrm{hom}}\to\operatorname{H}^{1}({\rm Gal}_{\rm\mathbb{Q}},V).$ Since $h^{*}\tilde{\Delta}_{GS}=6\Delta_{GS,\infty}$, we conclude that ${\rm AJ}_{\ell}(\tilde{\Delta}_{GS})$ is also nonzero and torsion. On the other hand, we may use Magma to compute the characteristic polynomial of a Frobenius element acting on $\operatorname{H}^{1}(C_{{\rm\overline{\mathbb{Q}}}},{\rm\mathbb{Z}}_{7}(1))$, and hence on $V$ as well (via Lemma 2.7). We check that the action of the geometric Frobenius $\mathrm{Fr}_{11}$ at $11$ on $V$ satisfies
@@
Thus $\operatorname{H}^{1}({\rm Gal}_{\rm\mathbb{Q}},V)$ is torsion-free by [48, 6.1], which is a contradiction. ∎
@@
The Ceresa cycle $\kappa_{\infty}(C)$ of the curve $C\colon y^{3}=x^{4}+x^{2}+1$ has infinite order.